Calculate Labour force, employment rate, unemployment rate using
Labour force participation rate : 65,8%
working age population: 5 485 000
absorption rate : 48 %
unemployment rate (previous year) : 25%
1
Expert's answer
2019-10-22T09:15:45-0400

labor force is the number of adult population seeking or in employment


"\frac{65.8}{100}\times {5485000}=3609130"

Since absorption rate is 48%,unemployment reduced by


"0.48\times0.25=0.12"

New unemployment rate is


"0.25-0.12=0.13"


13%

while employment rate is 87%